# Dense Hopfield Networks in the Teacher-Student Setting

### v1

@@ Abstract
Dense Hopfield networks are known for their feature to prototype transition and adversarial robustness. However, previous theoretical studies have been mostly concerned with their storage capacity. We bridge this gap by studying the phase diagram of $p$-body Hopfield networks in the teacher-student setting of an unsupervised learning problem, uncovering ferromagnetic phases reminiscent of the prototype and feature learning regimes. On the Nishimori line, we find the critical size of the training set necessary for efficient pattern retrieval. Interestingly, we find that that the paramagnetic to ferromagnetic transition of the teacher-student setting coincides with the paramagnetic to spin-glass transition of the direct model, i.e. with random patterns. Outside of the Nishimori line, we investigate the learning performance in relation to the inference temperature and dataset noise. Moreover, we show that using a larger $p$ for the student than the teacher gives the student an extensive tolerance to noise. We then derive a closed-form expression measuring the adversarial robustness of such a student at zero temperature, corroborating the positive correlation between number of parameters and robustness observed in large neural networks. We also use our model to clarify why the prototype phase of modern Hopfield networks is adversarially robust.
@@ 1 Introduction
Hopfield networks are artificial neural networks that model associative memory [1]. In the Hopfield model, examples $\sigma\in\left\{-1,1\right\}^{N}$ of memories $\xi^{\mu}\in\left\{-1,1\right\}^{N}$, $\mu=1,\ldots,M$, are retrieved by sampling the Gibbs distribution of a $2$-body Hamiltonian $H\left[\sigma|\xi\right]$ at a given temperature $T$ [2]. Hopfield networks can be trained in a biologically plausible way using Hebb’s rule [1, 3], which leads to $H\left[\sigma|\xi\right]=-\frac{1}{N}\sum_{\mu=1}^{M}\left(\sum_{i=1}^{N}\xi^{\mu}_{i}\sigma_{i}\right)^{2}$. However, they can only store up to $M\sim\mathcal{O}\left(N\right)$ random memories in the limit of large $N$ [1, 4, 5]. One way to find this scaling is to study the phase diagram of $H\left[\sigma|\xi\right]$ as a function of the temperature $T$ and load $\alpha=\frac{M}{N}$ [5], where the so-called ferromagnetic phase, which extends up to $\alpha\approx 0.14$, corresponds to accurate retrieval.
@@
Since Hopfield’s seminal work, several generalizations have been investigated in relation to their critical storage capacity and retrieval capabilities. For example, parallel retrieval has been studied in relation to pattern sparsity [6, 7, 8, 9, 10] or hierarchical interactions [11, 12, 13, 14, 15], and non-universality has been shown with respect to more general pattern entries and unit priors [16, 17, 18, 19, 20, 21, 22]. Efforts to overcome the $\mathcal{O}\left(N\right)$ limitation of the capacity led to the development of a novel class of modern Hopfield networks [23, 24, 25], which are sometimes called dense due to their faculty to store much more memories than the original Hopfield model [26]. These neural networks surpass $\mathcal{O}\left(N\right)$ storage capacity by using higher-order interactions instead of the original $2$-body couplings [27, 28, 29, 30, 31, 32]. In particular, Gardner [30] calculated the replica-symmetric (RS) phase diagram of the Hamiltonian $H\left[\sigma|\xi\right]=-\sum_{i_{1}<...<i_{p}=1}^{N}J_{i_{1}...i_{p}}\sigma_{i_{1}}...\sigma_{i_{p}}$ with $p$-body interactions $J_{i_{1}...i_{p}}=\frac{p!}{N^{p-1}}\sum_{\mu=1}^{M}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}$ conditioned on random memories $\xi^{\mu}\in\left\{-1,1\right\}^{N}$, finding a $M=\mathcal{O}\left(N^{p-1}\right)$ storage capacity. These calculations were later extended to include the effects of one-step replica symmetry breaking (1RSB) [33].
@@
Although they draw a rather detailed picture of the retrieval of individual random memories, these results are not the end of the story. First of all, there is still significant uncertainty on the location of the paramagnetic to spin-glass phase transition because of numerical instability. Second of all, dense Hopfield networks have been rapidly gaining a renewed attention for reasons other than their storage capacity since Krotov’s recent paper [26], where they were used as a trainable machine learning architecture. For instance, they have been related to transformers [23, 34] and diffusion models [35, 36], and they were found to be significantly more explainable and adversarially robust than feedforward neural networks with ReLU activation functions [26, 37].
@@
At finite load $\alpha=\frac{M}{N}$, two regimes of pattern retrieval were found: example retrieval ($eR$) and signal retrieval ($sR$). In the $eR$ phase, the student tries to reconstruct $\xi^{*}$ by directly retrieving the examples $\sigma^{a}$, which is a good strategy provided that they are strongly correlated with $\xi^{*}$. In the $sR$ phase, on the other hand, retrieval is done by extracting subtle cues from weakly correlated examples. The two types of examples used in these two retrieval strategies are respectively called prototypes and features of $\xi^{*}$ [26]. Interestingly, a prototype regime and a feature regime were also observed by Krotov in dense Hopfield networks trained to classify real data [26], where it was found that the prototype regime is significantly more adversarially robust than the feature regime. In other words, the prototype regime is more resistant than the feature regime to small data perturbations that are specifically designed to cause incorrect classification [40, 41]. This prototype approach is arguably a big step toward designing adversarially robust neural networks, a long-standing problem that still lacks a fully satisfying solution [42, 43, 44].
In this work, we study the performance of $p$-body Hopfield networks in the teacher-student setting, revealing a prototype regime and a feature regime as in the $2$-body model. In Section 2, we review Gardner’s main results in studying $p$-body Hopfield models. In Section 3, we compute the phase diagram of these $p$-body models in the teacher-student setting. In Section 4.1, we discuss the transition to the retrieval phase in the inverse problem and compare it against the transition to the spin-glass phase in the direct problem. Despite their different nature, we show that these two transitions are equivalent on the Nishimori line where the teacher and the student have the same $p$ and $T$ [45, 46, 47, 48]. In Section 4.2, we discuss the phase diagram on the Nishimori line in more details. In Section 4.3 and Section 4.4, we discuss the phase diagram outside of the Nishimori line. First of all, we investigate the effect of using an inference temperature different from the dataset noise. Second of all, we reveal that using a larger $p$ for the student than the teacher gives the student an extensive tolerance against both teacher noise and pattern interference. Finally, in Section 4.5, we derive a closed-form expression that measures the adversarial robustness of the student at zero temperature and explain what our results reveal about the nature of adversarial attacks.
@@ 2 Overview of Gardner’s results
$H\left[\sigma|\xi\right]=-\sum_{i_{1}<...<i_{p}=1}^{N}J_{i_{1}...i_{p}}\sigma_{i_{1}}...\sigma_{i_{p}}=-\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}=1}^{N}\sum_{\mu=1}^{M}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma_{i_{1}}...\sigma_{i_{p}}$ (1)
@@
$r$ $=pq^{p-1}$
$k$ $=pm^{p-1},$
and the order parameters $m$ and $q$ are to be interpreted as expected overlaps. To be more precise, $m$ can be shown to be the expected overlap of a retrieval attempt $\sigma$ against one memory in the thermodynamic limit, i.e. $m=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma_{i}\right\rangle_{\xi,\sigma}$. Similarly, $q$ is the expected overlap between two retrieval attempts $\sigma^{1}$ and $\sigma^{2}$, i.e. $q=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\sigma^{1}_{i}\sigma^{2}_{i}\right\rangle_{\xi,\sigma}=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\left\langle\sigma_{i}\right\rangle^{2}_{\sigma}\right\rangle_{\xi}$. Intuitively, $q$ measures the tendency of the system to stay frozen in specific configurations rather than visiting all possible values of $\sigma$.
The resulting RS phase diagram (see Fig. 1) can be derived from the value of the order parameters as a function of three hyperparameters: the interaction order $p$, temperature $T=1/\beta$ and load $\alpha=\frac{Mp!}{N^{p-1}}$. There are four different phases:
@@
Figure 1: RS phase diagrams of the direct models with $p=3$ on the left and $p=10$ on the right. Accurate pattern retrieval is not possible in the paramagnetic phase ($P$) or in the spin-glass phase ($SG$), but it is possible in the local retrieval phase ($lR$) and in the global retrieval phase ($gR$). The ferromagnetic fixed point corresponding to accurate pattern retrieval is globally stable in the $gR$ phase, but locally stable in the $lR$ phase. The phase diagrams are inexact below the white dashed line where the total entropy of the paramagnetic phase becomes negative.
Gardner also calculated the exact $p\rightarrow\infty$ phase diagram without making any assumptions about replica symmetry [30]. The resulting paramagnetic to spin-glass ($P$-$SG$) phase transition line, given by $\beta^{2}\alpha=2\log 2$, coincides with the boundary of the region where the total entropy of the paramagnetic phase becomes negative. Therefore, the total entropy is of the system is always positive, as expected. Conversely, the finite $p$ phase diagrams obtained under the RS approximation have a limited scope because RSB must be taken into account in the $\beta^{2}\alpha>2\log 2$ region where the RS approximation of the total entropy becomes negative [30, 50]. Although this phase diagram was recently extended by including the effects of 1RSB [33], there is still significant uncertainty on the location of the large-$p$ $P$-$SG$ transition because of numerical instability issues. In fact, the finite-$p$ transition evaluated in [33] does not seem to get closer to the exact $p\rightarrow\infty$ transition as $p$ gets larger.
@@ 3 Teacher-student setting
×$\quad\times\log\sum_{\xi}\exp\left(\beta\frac{p!}{N^{p-1}}\sum_{a=1}^{M}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\right)$ (2)
in the thermodynamic limit $N\to\infty$. We then draw phase diagrams of the inverse problem as a function of $p^{*}$, $T^{*}=1/\beta^{*}$, $p$, $T=1/\beta$ and $\alpha$.
We first consider the case where $p^{*}=p$ and the only possible mismatch between the teacher and student networks is in the inverse temperature, i.e. $\beta^{*}\neq\beta$. At low $T^{*}$, the student’s task is easy. In fact, below the critical temperature $T_{\text{crit}}$ of the direct problem with one pattern (see Fig. 1), the teacher produces examples $\sigma^{a}$ that cluster around $\xi^{*}$. Therefore, the student can infer $\xi^{*}$ by aligning its pattern $\xi$ with the examples $\sigma^{a}$. This retrieval strategy works even when using a very small amount of examples (see [38]). Since the size of our dataset is extensive, the retrieval accuracy is maximum in the thermodynamic limit. We call this region the (accurate) example Retrieval phase ($eR$).
Conversely, when $T^{*}$ is above $T_{\text{crit}}$, the examples in the training set are very noisy and we do not observe a finite overlap between $\sigma^{a}$ and $\xi^{*}$. In this regime, we find that the RS approximation of the free entropy can be computed (see Appendix D) in terms of the variational principle
$f$ $*\displaystyle=\operatorname*{Extr}_{m,k,q,r,q^{*},r^{*}}\Bigg{\{}\beta^{*}\beta\alpha\left[q^{*}\right]^{p}-\frac{1}{2}\beta^{2}\alpha q^{p}+\beta m^{p}-\beta^{*}\beta\alpha r^{*}q^{*}$ (3)

### v2

@@ Abstract
Dense Hopfield networks with $p$-body interactions are known for their feature to prototype transition and adversarial robustness. However, theoretical studies have been mostly concerned with their storage capacity. We derive the phase diagram of pattern retrieval in the teacher-student setting of $p$-body networks, finding ferromagnetic phases reminiscent of the prototype and feature learning regimes. On the Nishimori line, we find the critical amount of data necessary for pattern retrieval, and we show that the corresponding ferromagnetic transition coincides with the paramagnetic to spin-glass transition of $p$-body networks with random memories. Outside of the Nishimori line, we find that the student can tolerate extensive noise when it has a larger $p$ than the teacher. We derive a formula for the adversarial robustness of such a student at zero temperature, corroborating the positive correlation between number of parameters and robustness in large neural networks. Our model also clarifies why the prototype phase of $p$-body networks is adversarially robust.
This work is a SciPost Physics paper.
DOI: 10.21468/SciPostPhys.17.2.040.
@@ 1 Introduction
Hopfield networks are artificial neural networks that model associative memory [1]. In the Hopfield model, examples $\sigma\in\left\{-1,1\right\}^{N}$ of memories $\xi^{\mu}\in\left\{-1,1\right\}^{N}$, $\mu=1,\ldots,M$, are retrieved by sampling the Gibbs distribution of a $2$-body Hamiltonian $H\left[\sigma|\xi\right]$ at a given temperature $T$ [2]. Hopfield networks can be trained in a biologically plausible way using Hebb’s rule [1, 3], which leads to $H\left[\sigma|\xi\right]=-\frac{1}{N}\sum_{\mu=1}^{M}\left(\sum_{i=1}^{N}\xi^{\mu}_{i}\sigma_{i}\right)^{2}$. However, they can only store up to $M\sim\mathcal{O}\left(N\right)$ i.i.d. random memories in the limit of large $N$ [1, 4, 5]. One way to find this scaling is to study the phase diagram of $H\left[\sigma|\xi\right]$ as a function of the temperature $T$ and load $\alpha=\frac{M}{N}$ [5], where the so-called ferromagnetic phase, which extends up to $\alpha\approx 0.14$, corresponds to accurate retrieval.
@@
Since Hopfield’s seminal work, several generalizations have been investigated in relation to their critical storage capacity and retrieval capabilities. For example, parallel retrieval has been studied in relation to pattern sparsity [6, 7, 8, 9, 10] or hierarchical interactions [11, 12, 13, 14, 15], and non-universality has been shown with respect to more general pattern entries and unit priors [16, 17, 18, 19, 20, 21, 22]. Efforts to overcome the $\mathcal{O}\left(N\right)$ limitation of the capacity led to the development of a novel class of modern Hopfield networks [23, 24, 25], which are sometimes called dense due to their faculty to store much more memories than the original Hopfield model [26]. These neural networks surpass $\mathcal{O}\left(N\right)$ storage capacity by using higher-order interactions instead of the original $2$-body couplings [27, 28, 29, 30, 31, 32]. In particular, Gardner [30] calculated the replica-symmetric (RS) phase diagram of the Hamiltonian $H\left[\sigma|\xi\right]=-\sum_{i_{1}<...<i_{p}=1}^{N}J_{i_{1}...i_{p}}\sigma_{i_{1}}...\sigma_{i_{p}}$ with $p$-body interactions $J_{i_{1}...i_{p}}=\frac{p!}{N^{p-1}}\sum_{\mu=1}^{M}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}$ conditioned on i.i.d. random memories $\xi^{\mu}\in\left\{-1,1\right\}^{N}$, finding a $M=\mathcal{O}\left(N^{p-1}\right)$ storage capacity. These calculations were later extended to include the effects of one-step replica symmetry breaking (1RSB) [33].
@@
Although they draw a rather detailed picture of the retrieval of individual i.i.d. random memories, these results are not the end of the story. First of all, 1RSB calculations allegedly struggle to find the paramagnetic to spin-glass phase transition accurately at large $p$ because of numerical instability issues [33]. Second of all, dense Hopfield networks have been rapidly gaining a renewed attention for reasons other than their storage capacity since a recent paper [26] by Krotov and Hopfield (K & H), where they were used as a trainable machine learning architecture. For instance, they have been related to transformers [23, 34] and diffusion models [35, 36], and they were found to be significantly more explainable and adversarially robust than feedforward neural networks with ReLU activation functions [26, 37].
@@
At finite load $\alpha=\frac{M}{N}$, two regimes of pattern retrieval were found: example retrieval ($eR$) and signal retrieval ($sR$). In the $eR$ phase, the student tries to reconstruct $\xi^{*}$ by directly retrieving the examples $\sigma^{a}$, which is a good strategy provided that they are strongly correlated with $\xi^{*}$. In the $sR$ phase, on the other hand, retrieval is done by extracting subtle cues from weakly correlated examples. The two types of examples used in these two retrieval strategies are respectively called prototypes and features of $\xi^{*}$ [26]. Interestingly, a prototype regime and a feature regime were also observed by K & H in dense Hopfield networks trained to classify real data [26], where it was found that the prototype regime is significantly more adversarially robust than the feature regime. In other words, the prototype regime is more resistant than the feature regime to small data perturbations that are specifically designed to cause incorrect classification [40, 41]. This prototype approach is arguably a big step towards designing adversarially robust neural networks, a long-standing problem that still lacks a fully satisfying solution [42, 43, 44].
In this work, we study the performance of $p$-body Hopfield networks in the teacher-student setting, revealing a prototype regime and a feature regime as in the $2$-body model. In Section 2, we review Gardner’s main results in studying $p$-body Hopfield models and summarize what the rest of the literature on spin-glass models with $p$-body interactions tell us about the paramagnetic to spin-glass phase transition in $p$-body Hopfield models. In Section 3, we compute the phase diagram of these $p$-body models in the teacher-student setting. In Section 4.1, we discuss the transition to the retrieval phase in the inverse problem. In Section 4.2, we compare this retrieval transition against the transition to the spin-glass phase in the direct problem. Despite their different nature, we show that these two transitions are equivalent on the Nishimori line where the teacher and the student have the same $p$ and $T$ [45, 46, 47, 48]. In Section 4.3, we discuss the phase diagram on the Nishimori line in more details. In Section 4.4 and Section 4.5, we discuss the phase diagram outside of the Nishimori line. First of all, we investigate the effect of using an inference temperature different from the dataset noise. Second of all, we reveal that using a larger $p$ for the student than the teacher gives the student an extensive tolerance against both teacher noise and pattern interference. Finally, in Section 4.6, we derive a closed-form expression that measures the adversarial robustness of the student at zero temperature and explain what our results reveal about the nature of adversarial attacks.
@@ 2 Overview of Gardner’s results
$H\left[\sigma|\xi\right]=-\sum_{i_{1}<...<i_{p}=1}^{N}J_{i_{1}...i_{p}}\sigma_{i_{1}}...\sigma_{i_{p}}=-\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}=1}^{N}\sum_{\mu=1}^{M}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma_{i_{1}}...\sigma_{i_{p}}\,,$ (1)
@@
$r$ $=pq^{p-1}\,,$
$k$ $=pm^{p-1}\,,$ (2)
and the order parameters $m$ and $q$ are to be interpreted as expected overlaps. To be more precise, $m$ can be shown to be the expected overlap of a retrieval attempt $\sigma$ against one memory in the thermodynamic limit, i.e. $m=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma_{i}\right\rangle_{\xi,\sigma}$. Similarly, $q$ is the expected overlap between two retrieval attempts $\sigma^{1}$ and $\sigma^{2}$, i.e. $q=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\sigma^{1}_{i}\sigma^{2}_{i}\right\rangle_{\xi,\sigma}$ or equivalently $q=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\left\langle\sigma_{i}\right\rangle^{2}_{\sigma}\right\rangle_{\xi}$. Intuitively, $q$ measures the tendency of the system to stay frozen in specific configurations rather than visiting all possible values of $\sigma$.
The resulting RS phase diagram (see Fig. 1) are derived from the value of the order parameters as a function of three hyperparameters: the interaction order $p$, temperature $T=1/\beta$ and load $\alpha=\frac{Mp!}{N^{p-1}}$. There are four different phases:
@@
Figure 1: RS phase diagrams of the direct models with $p=3$ on the left and $p=10$ on the right. Accurate pattern retrieval is not possible in the paramagnetic phase ($P$) or in the spin-glass phase ($SG$), but it is possible in the local retrieval phase ($lR$) and in the global retrieval phase ($gR$). The ferromagnetic fixed point corresponding to accurate pattern retrieval is globally stable in the $gR$ phase, but locally stable in the $lR$ phase. The phase diagrams are inexact below the white dashed line where the total entropy of the paramagnetic phase becomes negative. The black dotted line overlaying the $p=3$ diagram is the (exact) 1RSB $P$-$SG$ transition temperature $T_{s}\left(\alpha,3\right)$, which is obtained by rescaling by $\sqrt{2\alpha}$ the corresponding transition temperature of the spin-glass model with $p$-body Gaussian interactions. The d1RSB transition $T_{d}\left(\alpha,3\right)$ is very close to $T_{s}\left(\alpha,3\right)$ throughout the displayed range of $\alpha$. The white dotted line in the $p=3$ plot is the temperature $T_{G}\left(\alpha,3\right)$ below which multiple steps of RSB are required to compute the free entropy. It is also obtained by rescaling by $\sqrt{2\alpha}$ the corresponding transition temperature of the Gaussian spin-glass model.
Gardner also calculated the exact $p\rightarrow\infty$ phase diagram without making any assumptions about replica symmetry [30]. In this limit, the resulting paramagnetic to spin-glass ($P$-$SG$) phase transition occurs at a temperature $T_{E}(\alpha)$ that coincides with the boundary of the region where the total entropy of the paramagnetic phase becomes negative, given by $\beta^{2}\alpha=2\log 2$ (white dashed line in Fig. 1).
At finite $p$, Gardner’s results only tell us that the model cannot be in the paramagnetic phase below $T_{E}(\alpha)$. Therefore, a spin-glass transition should occur at a temperature $T_{s}\left(\alpha,p\right)\geq T_{E}(\alpha)$. Since the RS spin glass solution of Eqs. (2) exists only below $T_{E}(\alpha)$ (violet region in Fig. 1), the spin-glass transition must be towards a RSB spin-glass phase.
Outside of the signal retrieval phases, the free entropy of the direct model is the same as for the spin-glass model with $p$-body Gaussian interactions where the temperature is rescaled by a factor of $\sqrt{2\alpha}$ [50, 51]. Therefore, the spin-glass and paramagnetic solutions are the same in the direct model as in this Gaussian spin-glass model, and we expect the exact phase diagrams of both models to be identical when the direct model is not in its signal retrieval phases. According to previous work on the Gaussian model with finite $p$ [51], a 1RSB solution with $m=k=0$ exists and is globally stable throughout a whole phase below $T_{s}(\alpha,p)\geq T_{E}(\alpha)$ (see Fig. 1). This solution becomes unstable at a lower transition temperature $T_{G}(\alpha,p)$ (see Fig. 2), below which multiple steps of RSB are required. In the limit of $p\to\infty$, it holds that $T_{s}(\alpha,p)\to T_{E}(\alpha)$ and $T_{G}(\alpha,p)\to 0$. In other terms, the direct model becomes 1RSB, which is consistent with the fact that it is converging to a random energy model with temperature rescaled by $\sqrt{2\alpha}$ [52, 50, 30]. Finally, we mention that this type of models exhibits a random first order transition phenomenology [53, 54, 55, 56]: there is in fact a range of temperatures $T_{s}(\alpha,p)\leq T\leq T_{d}(\alpha,p)$ where the dynamics get trapped in an exponential number of metastable clusters, with an emerging RSB structure that does not affect the free energy (see Fig. 2). This range of temperatures thus defines a so-called dynamical 1RSB (d1RSB) phase. Below $T_{s}(\alpha,p)$, the number of clusters is no longer exponential, and the system undergoes the thermodynamic 1RSB phase transition that we mentioned previously. The critical temperatures $T_{G}(\alpha,p)$, $T_{s}(\alpha,p)$ and $T_{d}(\alpha,p)$ can all be obtained by standard RSB methods, but the resulting saddle-point equations can be prone to numerical instability at large $p$ [33]. In Sections 4.2 and 4.3, we discuss an alternative way to obtain $T_{s}(\alpha,p)$ and $T_{d}(\alpha,p)$.
@@ 3 Teacher-student setting
×$\quad\times\log\sum_{\xi}\exp\left(\beta\frac{p!}{N^{p-1}}\sum_{a=1}^{M}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\right)\,,$ (3)
in the thermodynamic limit $N\to\infty$. We then draw phase diagrams of the inverse problem as a function of $p^{*}$, $T^{*}=1/\beta^{*}$, $p$, $T=1/\beta$ and $\alpha$, where $\alpha$ is $M$ normalized to $\mathcal{O}\left(1\right)$. Unless explicitly specified otherwise, we use $\alpha=\frac{Mp!}{N^{p-1}}$.
3.1 Matched interaction orders
We first consider the case where $p^{*}=p$ and the only possible mismatch between the teacher and student networks is in the inverse temperature, i.e. $\beta^{*}\neq\beta$. At low $T^{*}$, the student’s task is easy. In fact, below the critical temperature $T_{\text{crit}}$ of the direct problem with one pattern (see Fig. 1, $\alpha=0$ axis), the teacher produces examples $\sigma^{a}$ that cluster around $\xi^{*}$. Therefore, the student can infer $\xi^{*}$ by aligning its pattern $\xi$ with the examples $\sigma^{a}$. This retrieval strategy works even when using a very small amount of examples (see [38]). Since the size of our dataset is extensive, the retrieval accuracy is maximum in the thermodynamic limit. We call this region the (accurate) example Retrieval phase ($eR$).
Conversely, when $T^{*}$ is above $T_{\text{crit}}$, the examples in the training set are very noisy and we do not observe a finite overlap between $\sigma^{a}$ and $\xi^{*}$ (see Fig. 1, $\alpha=0$ axis). In this regime, we find that the RS approximation of the $p^{*}=p$ free entropy can be computed (see Appendix D) in terms of the variational principle
$f$ $=\operatorname*{Extr}_{m,k,q,r,q^{*},r^{*}}\Bigg{\{}\beta^{*}\beta\alpha\left[q^{*}\right]^{p}-\frac{1}{2}\beta^{2}\alpha q^{p}+\beta m^{p}-\beta^{*}\beta\alpha r^{*}q^{*}$